OpenStudy (anonymous):

The perimeter of a rectangle is 156m. The length 159m more than twice the width. find the dimensions.

OpenStudy (anonymous):

the whole perimeter is 156 and the length alone is more than 159m? are you sure the question is right?

OpenStudy (anonymous):

Oops sorry the length is 9m more than twice the width. my bad been up all night doing algebra.

OpenStudy (anonymous):

l=2b+9 perimeter=2(l+b) 156=2(2b+9+b) 156=2(3b+9) 78=3b+9 69=3b 23=b l=2*23 + 9=46+9=55